OpenStudy (anonymous):

The probability that it will rain today is 3/8. What is the propability that it will not rain today?

jimthompson5910 (jim_thompson5910):

P(Rain) + P(Not Rain) = 1 P(Not Rain) = 1 - P(Rain)

OpenStudy (anonymous):

im confused

jimthompson5910 (jim_thompson5910):

the probability that it rains + probability that it doesn't rain = 1

jimthompson5910 (jim_thompson5910):

since those are the only two options

jimthompson5910 (jim_thompson5910):

so we represent it like what I wrote above

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

to find the answer, simply plug in P(Rain) = 3/8 to get P(Not Rain) = 1 - P(Rain) P(Not Rain) = 1 - 3/8

OpenStudy (anonymous):

okay! so cross multiply?

jimthompson5910 (jim_thompson5910):

close, but not really

jimthompson5910 (jim_thompson5910):

1 = 8/8, so P(Not Rain) = 1 - 3/8 P(Not Rain) = 8/8 - 3/8 Now subtract

OpenStudy (anonymous):

So 1/4th?

jimthompson5910 (jim_thompson5910):

8 - 3 = ???

OpenStudy (anonymous):

@irishmolly not to butt in but i think you are thinking way to hard suppose it rains three days out of 8, how many days out of 8 does it not rain?

OpenStudy (anonymous):

*too hard

OpenStudy (anonymous):

4 days???

OpenStudy (anonymous):

thats my problem i overthink alot of these problems

OpenStudy (anonymous):

three out of eight people in the class are male how many are female?

OpenStudy (anonymous):

five females

OpenStudy (anonymous):

Boom! Got it :)

OpenStudy (anonymous):

three out of eight days it rains how many days does it not rain?

OpenStudy (anonymous):

five days so 5/8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much for helping me!!!

OpenStudy (anonymous):

yw